# Mechanical Sensors for Ultraheavy Dark Matter Searches via Long-range Forces

### v3

@@ II Detector Concept
where $N_{\mathrm{nuclei}}$ is the number of nuclei in the test mass and $r$ is the distance between the dark matter particle and the test mass. We parameterize the interaction strength based on nuclear number instead of nucleon number, $A$, because the typically assumed $A^{4}$ scaling for dark matter interaction cross section is not always valid for ultraheavy dark matter. Specifically, the assumptions required for $A^{4}$ scaling do not hold for large cross sections and non-contact interactions, and become model dependent [35]. While we study the detection of dark matter that couples to our sensor array via non-contact interactions, we also note that attractive forces interacting via a light mediator of mass $m_{\phi}$ would be indistinguishable from a long-range force as long as $1/m_{\phi}\gg d$, where $d$ is the sensor spacing.
@@ II.1.1 Statistical Track-finding
One downside of this approach is that doing a likelihood scan in high dimensions is computationally expensive, as the number of likelihood computations is $N^{d}$ for a $d$-dimensional parameter space with $N$ bins per parameter. An alternative approach is to use nested sampling, where the scaling of computational cost with dimensionality depends on the information gain when the number of parameters is increased, but can be as low as $O(d^{2})$ in some cases [40].
@@ IV.2 Sensitivity Projections
The near-term magnetic levitation temperature and mechanical quality factor $(Q_{\mathrm{m}})$ are based on [45], while the mass is increased significantly to increase the sensitivity. The mechanical quality factor of the near-term MEMS setup is chosen to be $10^{7}$, as macroscopic mechanical oscillators with similar or higher quality factors have been demonstrated in mm-scale high-stress silicon nitride membranes [61]; as such, we believe this is achievable for our application as long as tension is successfully retained in the silicon nitride spring. The temperature is chosen to be achievable with dilution refrigeration technology in large cm to metre scale volumes [62, 63]. Limitations to the laser power due to heating is also considered for the near-term MEMS setup, based on a cooling power of $200\,\mathrm{\mu W}$, achievable with state of the art dilution refrigerators [63] (see Appendix B). The near-term MEMS setup is given $10\,\mathrm{dB}$ of quantum noise reduction. This near-term quantum noise reduction for the optically accessible MEMS devices is based on the achieved 9 dB of squeezing demonstrated with rubidium vapor cell based free-space squeezed light sources [64, 65]. In addition, the 15 dB of squeezing used in the future milestone is based on the record squeezing that has been achieved with any source [66]. Finally, we take vibrational noise to be subdominant; this has been achieved at similar temperatures around $1\,\mathrm{Hz}$ [67]. The sensor material density is based on silicon nitride for the near-term MEMS setup as it is a favored material for high stress MEMS devices [61], and lead is used for the others as it is a dense type-I superconductor that has been demonstrated for chip-based magnetic levitation [46].
@@ Appendix A Conversion between cross section and coupling strength
where $\sigma$ is the per-nucleus cross section, $E_{r}^{\text{min}}$ is the minimum recoil energy, $E_{r}^{\text{max}}$ is the maximum recoil energy, $\alpha_{\mathrm{B-L}}$ is the BLB-L coupling strength that is analogous to the fine-structure constant, $A$ is the atomic mass number, $Z$ is the atomic number, $m_{n}$ is the mass of the target nucleus, $E_{\chi}$ is the energy of the incoming dark matter particle, and $m_{\chi}$ is the mass of the incoming dark matter particle.

### v4

@@ II Detector Concept
where $N_{\mathrm{nuclei}}$ is the number of nuclei in the test mass and $r$ is the distance between the dark matter particle and the test mass. We parameterize the interaction strength based on nuclear number instead of nucleon number, AA, because the typically assumed $A^{4}$ scaling for dark matter interaction cross section is not always valid for ultraheavy dark matter. Specifically, the assumptions required for $A^{4}$ scaling do not hold for large cross sections and non-contact interactions, and become model dependent [35]. While we study the detection of dark matter that couples to our sensor array via non-contact interactions, we also note that attractive forces interacting via a light mediator of mass $m_{\phi}$ would be indistinguishable from a long-range force as long as $1/m_{\phi}\gg d$, where $d$ is the sensor spacing.
@@ II.1.1 Statistical Track-finding
One downside of this approach is that doing a likelihood scan in high dimensions is computationally expensive, as the number of likelihood computations is $N^{d}$ for a $d$-dimensional parameter space with NN bins per parameter. An alternative approach is to use nested sampling, where the scaling of computational cost with dimensionality depends on the information gain when the number of parameters is increased, but can be as low as $O(d^{2})$ in some cases [40].
@@ IV.2 Sensitivity Projections
The near-term magnetic levitation temperature and mechanical quality factor $(Q_{\mathrm{m}})$ are based on [45], while the mass is increased significantly to increase the sensitivity. The mechanical quality factor of the near-term MEMS setup is chosen to be $10^{7}$, as macroscopic mechanical oscillators with similar or higher quality factors have been demonstrated in $m$-scale high-stress silicon nitride membranes [61]; as such, we believe this is achievable for our application as long as tension is successfully retained in the silicon nitride spring. The temperature is chosen to be achievable with dilution refrigeration technology in large cm to metre scale volumes [62, 63]. Limitations to the laser power due to heating is also considered for the near-term MEMS setup, based on a cooling power of $200\,\mathrm{\mu W}$, achievable with state of the art dilution refrigerators [63] (see Appendix B). The near-term MEMS setup is given $10\,\mathrm{dB}$ of quantum noise reduction. This near-term quantum noise reduction for the optically accessible MEMS devices is based on the achieved 9 dB of squeezing demonstrated with rubidium vapor cell based free-space squeezed light sources [64, 65]. In addition, the 15 dB of squeezing used in the future milestone is based on the record squeezing that has been achieved with any source [66]. Finally, we take vibrational noise to be subdominant; this has been achieved at similar temperatures around $1\,\mathrm{Hz}$ [67]. The sensor material density is based on silicon nitride for the near-term MEMS setup as it is a favored material for high stress MEMS devices [61], and lead is used for the others as it is a dense type-I superconductor that has been demonstrated for chip-based magnetic levitation [46].
@@ Appendix A Conversion between cross section and coupling strength
where $\sigma$ is the per-nucleus cross section, $E_{r}^{\text{min}}$ is the minimum recoil energy, $E_{r}^{\text{max}}$ is the maximum recoil energy, $\alpha_{\mathrm{B-L}}$ is the BLB-L coupling strength that is analogous to the fine-structure constant, AA is the atomic mass number, $Z$ is the atomic number, $m_{n}$ is the mass of the target nucleus, $E_{\chi}$ is the energy of the incoming dark matter particle, and $m_{\chi}$ is the mass of the incoming dark matter particle.